# Whitened Score Diffusion: A Structured Prior for Imaging Inverse Problems

### v3

@@ 3.1 Whitened Score matching objective
To mitigate these numerical instabilities in the score computation during training, we apply a whitening transformation to the score by naturally multiplying it with $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$, where $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\propto\boldsymbol{\Sigma}_{t}$, the forward diffusion process covariance. Similar to DSM, for our SDE in Eq. 8, we approximate $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ as $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ which has the following closed form expression after canceling $\boldsymbol{\Sigma}_{t}$:
@@ 3.2 Interpretation of WS
Concurrently with [4], [8] derived identical results for the reverse-time SDE, Eq. 2a, by decomposing the diffusion term of a reverse-time SDE into a unique sum of a zero-mean martingale and a predictable process $\mathbf{n}_{t}$, given as
@@ 3.3 WS diffusion priors for imaging inverse problems
We solve the imaging inverse problem using Eq. 7 with our WS diffusion prior and an approximation of the measurement likelihood score. Recall from Section 2.3 the myriad of methods developed to approximate the measurement likelihood score, all of which follow the template
@@
The reverse-time SDE framework aids inverse problems with correlated noise as the diffusion matrix $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ preconditions the inverse measurement covariance in the likelihood score, when $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ is designed to be proportional to the covariance matrix
@@
In practice, a regularization term $\lambda_{t}$ is important to balance the generative prior with the data likelihood. For proof-of-concept, we experiment with the likelihood-guided sampling via posterior score approximation in [19] due to its functional simplicity $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})\approx\boldsymbol{\Sigma}_{\mathbf{y}}^{-1}\mathbf{A}^{H}(\mathbf{y}-\mathbf{A}\mathbf{x}_{t})$. We also use the deterministic sampler of Eq. 7b. The resulting algorithm is shown in Algorithm 1.
@@ 4.1 Training details
For each dataset, we train two attention UNet models based on the architecture in [16] with three residual blocks in each downsampling layer, where one is for the conventional isotropic Gaussian SDE, and the other our anisotropic Gaussian SDE. We set the learning rate to $3\mathrm{e}^{-5}$ with a linear decay schedule. For CIFAR10 ($32\times 32$), the batch size is 128, and for CelebA-HQ ($64\times 64$), the batch size is 16. Models were trained on a single NVIDIA L40S GPU with 48GB of memory for two days. Our model is trained on the training sets of CIFAR-10 [25] and CelebA ($64\times 64$) [32] where $\mathbf{K}$ is a 2D Gaussian convolutional matrix characterized by an $\mathrm{std}$. For CIFAR, the $\mathrm{std}$ that characterizes $\mathbf{K}$ is uniformly distributed between 0.1 and 3, and between 0.1 and 5 for CelebA where $\mathrm{std}\leq 0.5$ equals the 2D delta function. The noise is also randomly grayscale or color with a 0.5 probability.
@@ 4.2 Imaging inverse problems with correlated noise
It is well established that natural image spectra exhibit exponential decay [47], indicating the dominance of low-frequency components in representing images. When additive measurement noise occupies the same spectral subspace, especially at low frequencies, the computational imaging task becomes fundamentally more challenging. We show that our framework is beneficial as a generative prior for solving inverse problems with such structured noise by experimenting with a variety of computational imaging modalities that are known to be affected by structured noise.
The measurements in our experiments are corrupted with additive grayscale structured noise, designed to mimic real-world conditions frequently encountered in both computational photography—such as fog, haze, and atmospheric turbulence—and computational microscopy—including fluorescence background, laser speckle, and detector noise. We use Algorithm 1 with $T=1000$ and $\beta_{min}=0.01$ and $\beta_{max}=20$ so that the $\mathrm{SNR}$ decays to 0 at $t=T$. For results with our WS prior, $\mathbf{x}_{T}$ was drawn from $\mathcal{N}(\mathbf{0},\mathbf{KK}^{\top})$ with $\mathrm{std}=3$ and $\mathrm{std}=6$ for CIFAR and CelebA, respectively with grayscale color (all color channels are the same value). For conventional DM prior results $\mathbf{x}_{T}$ was drawn from $\mathcal{N}(\mathbf{0},\mathbf{I})$. All evaluation was performed on unseen validation dataset sample images picked uniformly randomly. The regularization parameter $\lambda$ scales the magnitude of the likelihood step to be proportional to the magnitude of the prior step as was done by [19]. Line search was used to find an optimal $\lambda$ that yielded a reconstruction with the highest $\mathrm{PSNR}$, where $\mathrm{PSNR}$ is defined as $\mathrm{PSNR}=20\cdot\log_{10}\left(\frac{1}{\mathrm{MSE}}\right)$ where $\mathrm{MSE}$ is the mean squared error between the reconstruction and the ground truth.
Our results are demonstrated on a variety of computational imaging task such as imaging through fog, motion deblurring, lens deblurring, linear inverse scattering, and differential defocus. More details are in Appendix C.
@@
Computational imaging  Using WS diffusion prior to solve inverse problems with non-identity forward operators outperforms traditional score-based diffusion priors in $\mathrm{PSNR}$. Noticeably, our diffusion prior is able to maintain fidelity to the color distribution for restoring measurements corrupted by grayscale fog-like noise, while conventional score-based diffusion priors fail to remove the noise, as seen in Figs. 5 and 7 for deblurring inverse problems. More results on other imaging inverse problems and on the CelebA dataset can be found in Appendix C and Figs. 8, 9, and 10.
@@ Appendix A Denoising Whitened Score matching
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\mathbf{z}$ (18)
@@
[…]
@@
Let $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ denote the Gaussian probability transition kernel associated with the forward-time SDE in Eq. 8. For a linear SDE in $\mathbf{x}_{t}$, the covariance $\boldsymbol{\Sigma}_{t}$ of the transition kernel is a scalar multiple of twice the diffusion matrix, $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ ([44]) provided the initial conditions of $\boldsymbol{\mu}(0)=\mathbf{x}_{0}$ and $\boldsymbol{\Sigma}(0)=\mathbf{0}$ for $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$:
$\boldsymbol{\Sigma}_{t}=c\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ (21)
@@
$\min_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0}),\mathbf{x}_{0}\sim p(\mathbf{x})}\left[\|\mathbf{h}_{\boldsymbol{\theta}}(\mathbf{x}_{t})-\mathbf{x}_{0}\|_{2}^{2}\right]$ (22)
such that $\mathbf{h}_{\boldsymbol{\theta}^{*}}(\mathbf{x}_{t})=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$, for optimal network parameters $\boldsymbol{\theta}^{*}$. Tweedie’s formula from Lemma A.1 gives us that
@@
$\mathbf{h}_{\boldsymbol{\theta}}(\mathbf{x}_{t})=\frac{\mathbf{x}_{t}+c\mathbf{n}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{\alpha_{t}}$ (24)
@@
$\min_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0}),\mathbf{x}_{0}\sim p(\mathbf{x})}\left[\|\mathbf{x}_{t}+c\mathbf{n}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\alpha_{t}\mathbf{x}_{0}\|_{2}^{2}\right]$ (25)
@@
which proves that our model $\mathbf{n}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ learns $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ with objective Eq. 20
@@ Appendix B Flow matching in SDE
$p_{t}=\mathcal{N}(\mathbf{x}_{t}\mid\boldsymbol{\mu}_{t}(\mathbf{x}_{0}),\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0}))$ (27)
@@
$\phi_{t}(\mathbf{x}_{0})=\boldsymbol{\mu}_{t}(\mathbf{x}_{0})+\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0})$ (28)
where we define $\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0})$ such that $Cov[\phi_{t}(\mathbf{x}_{0})]=\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0})=\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0})(\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0}))^{\top}$ and the initial condition $\boldsymbol{\mu}_{0}(\mathbf{x}_{0})=\mathbf{x}_{0}$. This probability path is equivalent to the probability transition kernel in Eq. 9 defined by a linear SDE Eq. 8 with drift coefficient $\mathbf{F}_{t}$ and diffusion matrix $\mathbf{G}_{t}$. Therefore we may attain the time derivatives of the mean and covariance functions using Fokker-Planck (see Eqs. 6.2 in [44]) expressed as
$\boldsymbol{\Sigma}_{t}^{\prime}(\mathbf{x}_{0})=2\mathbf{F}_{t}\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0})+\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ (29)
@@
$=\mathbf{F}_{t}(2\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0})-\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ (35)

### v4

@@ 3.1 Whitened Score matching objective
To mitigate these numerical instabilities in the score computation during training, we apply a whitening transformation to the score by naturally multiplying it with $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$, where $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\propto\boldsymbol{\Sigma}_{t}$, the forward diffusion process covariance. Similar to DSM, for our SDE in Eq. 8, we approximate $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ as $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ which has the following closed-form expression after canceling $\boldsymbol{\Sigma}_{t}$ with $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$:
@@
This objective defines a new learning target within the broader landscape of diffusion model losses. Our approach can be seen as a generalization of noise prediction [ho_denoising_2020] to the setting of correlated Gaussian noise, where the preconditioning term $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ captures the noise structure. Unlike conventional noise prediction, which assumes isotropic noise, our formulation enables stable training under arbitrary Gaussian forward processes. Furthermore, Eq. 15 reveals that the conditional FM vector field $\mathbf{u}_{t}$ is a linear combination of our conditional WS function and the drift term (see Appendix B), highlighting that both WS and FM avoid covariance inversion by preconditioning the score function with $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$. This shared property enables principled modeling of flexible Gaussian probability paths.
@@ 3.2 Interpretation of WS
Concurrently with [anderson_reverse-time_1982], [castanon_reverse-time_1982] derived identical results for the reverse-time SDE, Eq. 2a, by decomposing the diffusion term of a reverse-time SDE into a unique sum of a zero-mean martingale and a predictable process $\mathbf{n}_{t}$, given as
@@ 3.3 WS diffusion priors for imaging inverse problems
We solve the imaging inverse problem using Eq. 7 with our WS diffusion prior and an approximation of the measurement likelihood score. Recall from Section 2.4 the myriad of methods developed to approximate the measurement likelihood score, all of which follow the template,
@@
The reverse-time SDE framework aids inverse problems with correlated noise as the diffusion matrix $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ preconditions the inverse measurement covariance in the likelihood score, when $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ is designed to be proportional to the covariance matrix,
@@
In practice, a regularization term $\lambda_{t}$ is important to balance the generative prior with the data likelihood. For proof-of-concept, we experiment with the likelihood-guided sampling via posterior score approximation in [jalal_robust_2021] due to its functional simplicity $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})\approx\boldsymbol{\Sigma}_{\mathbf{y}}^{-1}\mathbf{A}^{H}(\mathbf{y}-\mathbf{A}\mathbf{x}_{t})$. We also use the deterministic sampler of Eq. 7b. The resulting algorithm is shown in Algorithm 1.
@@ 4.1 Training details
For each dataset, we train two attention UNet models based on the architecture in [ho_denoising_2020] with three residual blocks in each downsampling layer, where one is for the conventional isotropic Gaussian SDE, and the other our anisotropic Gaussian SDE. We set the learning rate to $3\mathrm{e}^{-5}$ with a linear decay schedule. For CIFAR10 ($32\times 32$), the batch size is 128, for CelebA ($64\times 64$), the batch size is 16, and for CelebA-HQ ($256\times 256$), the batch size is 4. Models were trained on a single NVIDIA L40S GPU with 48GB of memory for two days. Our model is trained on the training sets of CIFAR-10 [krizhevsky_learning_nodate], CelebA ($64\times 64$) [liu2015faceattributes] and CelebA-HQ ($256\times 256$) [karras2017progressive] where $\mathbf{K}$ is a 2D Gaussian convolutional matrix characterized by an $\mathrm{std}$. For CIFAR, the $\mathrm{std}$ that characterizes $\mathbf{K}$ is uniformly distributed between 0.1 and 3, between 0.1 and 5 for CelebA ($64\times 64$), and between 0.1 and 20 for CelebA-HQ ($256\times 256$) where $\mathrm{std}\leq 0.5$ equals the 2D delta function. The noise is also randomly grayscale or color with a 0.5 probability.
@@ 4.2 Imaging inverse problems with correlated noise
It is well-established that natural image spectra exhibit exponential decay [VANDERSCHAAF19962759], indicating the dominance of low-frequency components in representing images. When additive measurement noise occupies the same spectral subspace, especially at low frequencies, the computational imaging task becomes fundamentally more challenging. We show that our framework is beneficial as a generative prior for solving inverse problems with such structured noise by experimenting with a variety of computational imaging modalities that are known to be affected by structured noise.
The measurements in our experiments are corrupted by additive grayscale structured noise, designed to mimic real-world conditions frequently encountered in both computational photography—such as fog, haze, and atmospheric turbulence—and computational microscopy—including fluorescence background, laser speckle, and detector noise. We use Algorithm 1 with $T=1000$ and $\beta_{min}=0.01$ and $\beta_{max}=20$ so that the $\mathrm{SNR}$ decays to 0 at $t=T$. For results with our WS prior, $\mathbf{x}_{T}$ was drawn from $\mathcal{N}(\mathbf{0},\mathbf{KK}^{\top})$ with $\mathrm{std}=3$ and $\mathrm{std}=6$ for CIFAR and CelebA, respectively with grayscale color (all color channels have the same value). For conventional DM prior results $\mathbf{x}_{T}$ was drawn from $\mathcal{N}(\mathbf{0},\mathbf{I})$. All evaluation was performed on unseen validation dataset sample images picked uniformly at random. The regularization parameter $\lambda$ scales the magnitude of the likelihood step to be proportional to the magnitude of the prior step as was done by [jalal_robust_2021]. Line search was used to find an optimal $\lambda$ that yielded a reconstruction with the highest $\mathrm{PSNR}$, where $\mathrm{PSNR}$ is defined as $\mathrm{PSNR}=20\cdot\log_{10}\left(\frac{1}{\mathrm{MSE}}\right)$ where $\mathrm{MSE}$ is the mean squared error between the reconstruction and the ground truth.
Our results are demonstrated on a variety of computational imaging tasks such as imaging through fog, motion deblurring, lens deblurring, linear inverse scattering, and differential defocus. More details are in Appendix C.
@@
Computational imaging  Using WS diffusion prior to solve inverse problems with non-identity forward operators outperforms traditional score-based diffusion priors in $\mathrm{PSNR}$. Noticeably, our diffusion prior is able to maintain fidelity to the color distribution for restoring measurements corrupted by grayscale fog-like noise, while conventional score-based diffusion priors fail to remove the noise, as seen in Figs. 5 and 8 for deblurring inverse problems. Additional results for other imaging inverse problems on CIFAR, as well as on the CelebA ($64\times 64$) and CelebA-HQ ($256\times 256$) datasets, are presented in Appendix C and and Figs. 7, 8, 9, 10, and 11.
@@ 5 Conclusion
Limitations and future work
A primary limitation of our approach lies in the computational cost of sampling. The current time discretization of the reverse-time SDE necessitates approximately 1000 denoising steps, which may be prohibitive for certain practical applications. Reducing the number of denoising steps through model distillation represents a promising direction for future work [salimans_progressive_2022, song_consistency_2023].
Another limitation concerns the absence of an explicit mechanism to estimate the measurement noise covariance, which directly influences the specification of the diffusion matrix $\mathbf{G}_{t}$. A natural extension of this framework would involve parameterizing and learning $\mathbf{G}_{t}$ jointly with the model parameters. Such an approach would allow the diffusion process to adaptively capture data-dependent or task-specific noise structures, thereby enhancing the model’s flexibility and representational capacity. This line of work connects to recent advances in vector-valued and multivariate diffusion models, which have demonstrated improved performance in scenarios characterized by complex or structured noise.
Finally, while our model exhibits strong performance as a denoising prior, additional research is needed for WS DMs to achieve competitive results in unconditional or conditional generation tasks. Promising directions include latent diffusion formulations and related techniques for improving generative efficiency and expressiveness [chung2024prompttuning, rout_beyond_2023, rombach_high-resolution_2022].
@@ Appendix A Denoising Whitened Score matching
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\mathbf{z},$ (18)
@@
$=\frac{1}{p_{t}(\mathbf{x}_{t})}\int p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})p_{0}(\mathbf{x}_{0})\,\mathrm{d}\mathbf{x}_{0}$ $\text{\small(use identity }\nabla f=f\nabla\log f\text{)}$
@@
Let $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ denote the Gaussian probability transition kernel associated with the forward-time SDE in Eq. 8. For a linear SDE in $\mathbf{x}_{t}$, the covariance $\boldsymbol{\Sigma}_{t}$ of the transition kernel is a scalar multiple of twice the diffusion matrix, $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$ ([sarkka_applied_2019]) provided the initial conditions of $\boldsymbol{\mu}(0)=\mathbf{x}_{0}$ and $\boldsymbol{\Sigma}(0)=\mathbf{0}$ for $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$:
$\boldsymbol{\Sigma}_{t}=c\mathbf{G}_{t}\mathbf{G}_{t}^{\top}.$ (21)
@@
$\min_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0}),\mathbf{x}_{0}\sim p(\mathbf{x})}\left[\|\mathbf{h}_{\boldsymbol{\theta}}(\mathbf{x}_{t})-\mathbf{x}_{0}\|_{2}^{2}\right],$ (22)
such that $\mathbf{h}_{\boldsymbol{\theta}^{*}}(\mathbf{x}_{t})=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$, for optimal network parameters $\boldsymbol{\theta}^{*}$. Tweedie’s formula from Lemma A.1 gives us that,
@@
$\mathbf{h}_{\boldsymbol{\theta}}(\mathbf{x}_{t})=\frac{\mathbf{x}_{t}+c\mathbf{n}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{\alpha_{t}},$ (24)
@@
$\min_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0}),\mathbf{x}_{0}\sim p(\mathbf{x})}\left[\|\mathbf{x}_{t}+c\mathbf{n}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\alpha_{t}\mathbf{x}_{0}\|_{2}^{2}\right],$ (25)
@@
which proves that our model $\mathbf{n}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ learns $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ with objective Eq. 20.
@@ Appendix B Flow matching in SDE
$p_{t}=\mathcal{N}(\mathbf{x}_{t}\mid\boldsymbol{\mu}_{t}(\mathbf{x}_{0}),\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0})),$ (27)
@@
$\phi_{t}(\mathbf{x}_{0})=\boldsymbol{\mu}_{t}(\mathbf{x}_{0})+\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0}),$ (28)
where we define $\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0})$ such that $\text{Cov}[\phi_{t}(\mathbf{x}_{0})]=\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0})=\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0})(\boldsymbol{\Sigma}_{t}^{\frac{1}{2}}(\mathbf{x}_{0}))^{\top}$ and the initial condition $\boldsymbol{\mu}_{0}(\mathbf{x}_{0})=\mathbf{x}_{0}$. This probability path is equivalent to the probability transition kernel in Eq. 9 defined by a linear SDE Eq. 8 with drift coefficient $\mathbf{F}_{t}$ and diffusion matrix $\mathbf{G}_{t}$. Therefore we may attain the time derivatives of the mean and covariance functions using Fokker-Planck (see Eqs. 6.2 in [sarkka_applied_2019]) expressed as
$\boldsymbol{\Sigma}_{t}^{\prime}(\mathbf{x}_{0})=2\mathbf{F}_{t}\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0})+\mathbf{G}_{t}\mathbf{G}_{t}^{\top},$ (29)
@@
$=\mathbf{F}_{t}(2\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0})-\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{x}_{0}).$ (35)